OpenStudy (anonymous):

(3/2)x - (1/4)x = -(1/3)x + 5/6

OpenStudy (anonymous):

are you solving for x?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

(3/2)x - (1/4)x = -(1/3)x + 5/6 (3/2)x - (1/4)x + (1/3)x= (5/6) *find common denominators* (18/12)x -(3/12)x + (4/12)x = (5/6) ((18-3+4)/12) x= (5/6) (19/12)x=(5/6) multiply by reciprocal (12/19)(19/12)x = (5/6)(12/19) x=10/19 hope this made sense!

OpenStudy (anonymous):

ok got you thanks so much ,, you really helped